# ExAct: A Video-Language Benchmark for Expert Action Analysis

### v1

@@ Abstract
We present ExAct, a new video-language benchmark for expert-level understanding of skilled physical human activities. Our new benchmark contains 3,521 expert-curated video question-answer pairs spanning 11 physical activities in 6 domains: Sports, Bike Repair, Cooking, Health, Music, and Dance. ExAct requires the correct answer to be selected from five carefully designed candidate options, thus necessitating a nuanced, fine-grained, expert-level understanding of physical human skills. Evaluating the recent state-of-the-art VLMs on ExAct reveals a substantial performance gap relative to human expert performance. Specifically, the best-performing GPT-4o model achieves only 44.70% accuracy, well below the 82.02% attained by trained human specialists/experts. We believe that ExAct will be beneficial for developing and evaluating VLMs capable of precise understanding of human skills in various physical and procedural domains. Dataset and code are available at https://texaser.github.io/exact_project_page/.
@@ 1 Introduction
Kinetics-700 [8]
HowTo100M [33]
UCF101 [45]
HMDB [21]
Moments in Time [34]
Hollywood [43]
ActivityNet-QA [51]
@@
Something-SomethingV2 [15]
FineGym [42]
Multisports [25]
TemporalBench [6]
@@
JIGSAWS [1]
Best [12]
FineDiving [50]
FP-Basket [4]
BASKET [36]
Aifit [13]
@@
VidDiffBench [5]
EgoExo-Fitness [26]
EgoExolearn [19]
Ego-Exo4D [16]
@@
Today, learning and perfecting a new physical skill requires a significant amount of time, practice, and often guidance from an expert coach/professional in that domain. Unfortunately, personalized coaching remains inaccessible to the majority due to prohibitively expensive costs and a lack of expert availability. Recent advances in AI have sparked interest in developing virtual AI assistants/coaches, particularly in the text domain [35, 2, 11]. However, text-based large language models (e.g., ChatGPT) are insufficient for learning new physical skills as standard LLMs typically lack a nuanced understanding of physical human activities/skills. In contrast, rapidly improving vision-language models (VLMs) could serve as valuable tools for various physical skill learning applications. In particular, by recording a video of a skill demonstration and feeding it to a VLM, people could receive detailed, actionable feedback similar to that of expert human coaches.
@@
Although recent progress has enabled modern VLMs to achieve impressive general image/video recognition capabilities [20, 46, 29, 3, 27, 9], existing studies reveal that such VLMs still struggle to understand fine-grained human activities [6, 36], particularly activities that require expert-level knowledge [5, 19, 26]. This is primarily due to the inadequate underlying visual representations that 1) do not capture expert-level knowledge needed to generate feedback for physical skill learning, and 2) the inability to recognize subtle details in skilled human actions.
@@
In addition to the limitations of existing VLMs, there is a notable lack of evaluation benchmarks tailored for expert-level understanding of skilled human activities. As shown in Table 1, most existing datasets focus on coarse activity recognition [8, 33, 45, 21, 34, 43, 51], which typically only require scene-level recognition rather than fine-grained understanding. While several fine-grained video recognition datasets exist [15, 42, 25, 6], the majority of them are aimed at generic (e.g., putting something into something) rather than expert-level action understanding of skilled human activities (e.g., keeping balance during a dance spin, playing the correct rhythm in a piano piece). Beyond action recognition, a number of video-based skill assessment benchmarks have been recently developed [1, 12, 50, 4, 36, 13]. Most of such skill assessment datasets focus on predicting scalar/categorical performance scores to quantify execution quality. Additionally, these datasets often lack open-ended language annotations, which can provide a rich and intuitive medium to convey skill-specific feedback [5]. Finally, a few recent skilled activity video-language datasets [5, 26, 19, 16] use experts to obtain free-form language descriptions akin to verbal coach feedback but do not provide rigorous evaluation benchmarks or tasks, making it difficult to assess how well modern video models can understand nuanced physical human skills.
@@
To address these issues, we introduce ExAct, a video-language benchmark consisting of 3,521 video–question–answer (VQA) pairs designed to evaluate expert-level understanding of skilled physical human actions (sample questions shown in Figure 1). ExAct covers 11 activities across 6 diverse physical domains: Sports (Basketball, Soccer, Bouldering), Bike Repair, Cooking, Health (COVID-19 Safety, CPR), Music (Guitar, Piano, Violin), and Dance, as shown in Figure 2. To construct ExAct, we build on the fine-grained expert commentaries derived from the Ego-Exo4D dataset [16]. The original Ego-Exo4D expert commentaries have several crucial limitations: 1) the original expert commentaries are unstructured and lengthy, 2) they contain automatic speech recognition (ASR) errors, and 3) include redundant or irrelevant content that is not directly tied to the observed actions. Furthermore, evaluating the quality of open-ended captions/descriptions in the original form of Ego-Exo4D expert commentaries is inherently difficult using standard language metrics such as CIDEr [47], BLEU [37], or ROUGE [28], which do not accurately reflect the correctness or relevance of the instructional feedback in the context of skilled physical activities. To address these challenges, we first apply a structured annotation pipeline that rewrites the original commentaries into concise, self-contained feedback commentaries. Then, our proposed ExAct benchmark evaluates expert action understanding as a multiple-choice question-answering task, which uses a well-defined metric of question-answering accuracy.
We conduct extensive empirical evaluation on several state-of-the-art VLMs, including GPT-4o [20], Gemini 1.5 Pro [46], LLaVA-Video [55], LLaVA-OneVision [22], Qwen2.5-VL [3], VideoLLaMA [53], InternVL2.5 [9], and PerceptionLM [10] on our new ExAct benchmark. Our results reveal that most modern VLMs achieve poor results on ExAct. In particular, GPT-4o, the best-performing model in our experiments, achieves only 44.70% accuracy. In comparison, we observed that non-expert humans can achieve around 61.86% accuracy, while domain experts achieve 82.02% accuracy. This substantial gap highlights the limitations of current VLMs in expert-level understanding of physical human skills. We hope that ExAct will serve as a rigorous evaluation benchmark for measuring expert-level understanding of skilled human actions, thus laying the foundation for AI systems that support enhanced human skill learning.
@@ 2 Related Work
Vision-Language Models (VLMs). Recent advances in Vision-Language Models (VLMs) have demonstrated impressive capabilities in understanding visual content. Models such as GPT-4o [20], Gemini 1.5 [46], LLaVA-OneVision [22], Qwen2.5-VL [3], VideoLLaMA [53], and InternVL2.5 [9] have achieved strong performance in tasks such as action recognition, video captioning, and visual question answering. Although these models show remarkable generalization, their outputs are often limited to high-level descriptions of the image/video content, rather than a detailed, fine-grained understanding of physical human actions and skills. More recent video-centric variants, such as LLaVA-Video [55] and VideoLLaMA [53], attempt to extend static image capabilities to temporal inputs. However, these models still struggle to understand the fine-grained nuances of complex physical human activities, which are essential for human skill understanding. Specifically, most existing VLMs cannot assess how well a task is performed, nor articulate the strengths of the execution and the areas requiring improvement. To address this gap, our proposed ExAct provides a rigorous benchmark to evaluate expert-level understanding of physical human skills.
@@
Multimodal and VQA Benchmarks. Diverse multimodal benchmarks have emerged to evaluate VLM performance, particularly in the video domain. ActivityNet-QA [51] assesses the temporal reasoning of activities within longer videos via question-answering. NExT-QA [49] focuses on compositional temporal reasoning. STAR [48] emphasizes situated reasoning. Benchmarks such as PerceptionTest [41] probe the perceptual abilities of VLMs in various modalities. TemporalBench [6] targets fine-grained temporal understanding. MV-Bench [24] assesses temporal comprehension across 20 challenging tasks. EgoSchema [31] focuses on egocentric human actions. MMWorld [17] aims to evaluate embodied agents in simulated environments. MLVU [57], Video-MMMU [18], MMVU [56], Video-MMLU [44], and Video-MME [14] aim to evaluate modern MLLMs for complex video question answering tasks. Finally, SEED-Bench [23] and MMBench [30] evaluate various multimodal abilities of MLLMs. However, most existing benchmarks focus on factual recall (e.g., “What color is the car?”), event-level understanding (e.g., “What action is being performed?”), or basic temporal reasoning (e.g., “What happened before this?”), and are not designed to capture the subtle nuances of skilled human actions. In contrast, our newly proposed ExAct benchmark focuses explicitly on expert-level analysis of physical human skills.
@@
Video-based Skill Assessment Benchmarks. Several recent works have focused on developing methods to assess human skills from video. Benchmarks such as MITDive [40], UNLV-Dive [39], MTL-AQA [38], and FineDiving [50] provide temporally segmented videos with action labels or action quality scoring for diving. FP-Basket [4] and BASKET [36]) focus on basketball, while LOGO [54]) provides human judgment scores for artistic swimming. Similar efforts also exist in other sports, including figure skating [40] and golf [32]. Furthermore, datasets such as JIGSAWS [1], BEST [12], and EgoExoLearn [19] focus on scenarios beyond sports, such as surgical tasks and daily activities. Most recently, the Ego-Exo4D [16] dataset introduces large-scale egocentric and exocentric video of skilled human activities. Ego-Exo4D includes spoken expert commentaries, offering a unique expert-level supervisory signal for understanding human skills. However, these expert commentaries are typically highly unstructured, noisy due to ASR errors, and often contain irrelevant information. Moreover, Ego-Exo4D does not provide a formal evaluation benchmark/task associated with such expert commentaries. In our work, we leverage such unstructured expert commentaries and construct a rigorous, expert-validated, and easy-to-evaluate ExAct benchmark enabling evaluation of expert-level understanding of physical human actions/skills.
@@ 3 ExAct Benchmark Construction
We construct ExAct using a four-stage pipeline. In stage I, we pre-process raw expert commentaries using GPT-4o, correcting errors and segmenting them into concise, self-contained feedback commentaries. In Stage II, we construct multiple-choice QA pairs, each consisting of one correct commentary and four distractors. In Stage III, we remove low-quality or biased samples through length filtering and blind-LLMs. Finally, in Stage IV, domain experts review each QA pair to ensure visual grounding and linguistic accuracy. Our benchmark construction pipeline is illustrated in Figure 3.
@@ 3.3 Stage III: Generated Question Answer Filtering
After constructing initial QA samples (Subsection 3.2), we apply several additional filtering steps to ensure high-quality and unbiased samples. Prior studies [7, 52] have shown that large language models (LLMs) can often exploit subtle statistical or stylistic patterns in multiple-choice answers to correctly identify the ground-truth option without relying on visual input. Such language-driven shortcuts, often referred to as language-related bias, pose a serious threat to the integrity of robust evaluation. To mitigate such biases, we adopt two filtering strategies:
1) Length similarity: Similar to human test-takers, LLMs may exploit surface-level cues such as answer length to eliminate implausible distractors. To mitigate this bias, we enforce a length similarity constraint: each distractor must be between 80% and 120% the length of the positive (i.e., correct) expert commentary, and the absolute word count difference must not exceed 8 words. QA samples that violate this constraint are excluded from the benchmark.
2) Blind-LLM filtering: Inspired by findings from TemporalBench [6], we observe that LLMs can sometimes identify the correct answer by detecting shared linguistic patterns, particularly when distractors are lightly edited variants of the ground truth. To avoid such language-driven biases, we present each QA sample consisting solely of the five textual options without any video to a set of state-of-the-art LLMs. Each model is prompted with the question: “Which expert commentary best matches the provided video?” If more than 20% of the models (i.e., exceeding the random chance of selecting the correct answer in a 5-way setup) select the positive expert commentary, we consider the sample susceptible to language-only bias and remove it from the dataset.

### v2

@@ Abstract
We present ExAct, a new video-language benchmark for expert-level understanding of skilled physical human activities. Our new benchmark contains 3,521 expert-curated video question-answer pairs spanning 11 physical activities in 6 domains: Sports, Bike Repair, Cooking, Health, Music, and Dance. ExAct requires the correct answer to be selected from five carefully designed candidate options, thus necessitating a nuanced, fine-grained, expert-level understanding of physical human skills. Evaluating the recent state-of-the-art VLMs on ExAct reveals a substantial performance gap relative to human expert performance. Specifically, the best-performing Gemini 2.5 Pro model achieves only 55.35% accuracy, well below the 82.02% attained by trained human experts. We believe that ExAct will be beneficial for developing and evaluating VLMs capable of precise understanding of human skills in various physical and procedural domains. Dataset and code are available at https://texaser.github.io/exact_project_page/.
@@ 1 Introduction
Kinetics-700 [carreira2019short]
HowTo100M [miech2019howto100m]
UCF101 [soomro2012ucf101]
HMDB [kuehne2011hmdb]
Moments in Time [monfort2019moments]
Hollywood [sigurdsson2016hollywood]
ActivityNet-QA [yu2019activitynet]
@@
Something-SomethingV2 [goyal2017something]
FineGym [shao2020finegym]
Multisports [li2021multisports]
TemporalBench [cai2024temporalbench]
@@
JIGSAWS [ahmidi2017dataset]
Best [Doughty_2019_CVPR]
FineDiving [xu2022finediving]
FP-Basket [bertasius2017baller]
BASKET [pan2025basket]
Aifit [fieraru2021aifit]
@@
VidDiffBench [burgess2025video]
EgoExo-Fitness [li2024egoexo]
EgoExolearn [huang2024egoexolearn]
Ego-Exo4D [grauman2024ego]
@@
Today, learning and perfecting a new physical skill requires a significant amount of time, practice, and often guidance from an expert coach/professional in that domain. Unfortunately, personalized coaching remains inaccessible to the majority due to prohibitively expensive costs and a lack of expert availability. Recent advances in AI have sparked interest in developing virtual AI assistants/coaches, particularly in the text domain [openai2024chatgpt, anthropic2023claude, team2024gemini]. However, text-based large language models (e.g., ChatGPT) are insufficient for learning new physical skills as standard LLMs typically lack a nuanced understanding of physical human activities/skills. In contrast, rapidly improving vision-language models (VLMs) could serve as valuable tools for various physical skill learning applications. In particular, by recording a video of a skill demonstration and feeding it to a VLM, people could receive detailed, actionable feedback similar to that of expert human coaches.
@@
Although recent progress has enabled modern VLMs to achieve impressive general image/video recognition capabilities [hurst2024gpt, team2024gemini, liu2023visual, bai2025qwen2, lin2023video, chen2024internvl], existing studies reveal that such VLMs still struggle to understand fine-grained human activities [cai2024temporalbench, pan2025basket], particularly activities that require expert-level knowledge [burgess2025video, huang2024egoexolearn, li2024egoexo]. This is primarily due to the inadequate underlying visual representations that 1) do not capture expert-level knowledge needed to generate feedback for physical skill learning, and 2) are unable to recognize subtle details in skilled human actions.
@@
In addition to the limitations of existing VLMs, there is a notable lack of evaluation benchmarks tailored for expert-level understanding of skilled human activities. As shown in Table 1, most existing datasets focus on coarse activity recognition [carreira2019short, miech2019howto100m, soomro2012ucf101, kuehne2011hmdb, monfort2019moments, sigurdsson2016hollywood, yu2019activitynet], which typically only require scene-level recognition rather than fine-grained understanding. While several fine-grained video recognition datasets exist [goyal2017something, shao2020finegym, li2021multisports, cai2024temporalbench], the majority of them are aimed at generic (e.g., putting something into something) rather than expert-level action understanding of skilled human activities (e.g., keeping balance during a dance spin, playing the correct rhythm in a piano piece). Beyond action recognition, a number of video-based skill assessment benchmarks have been recently developed [ahmidi2017dataset, Doughty_2019_CVPR, xu2022finediving, bertasius2017baller, pan2025basket, fieraru2021aifit]. Most of such skill assessment datasets focus on predicting scalar/categorical performance scores to quantify execution quality. Additionally, these datasets often lack open-ended language annotations, which can provide a rich and intuitive medium to convey skill-specific feedback [burgess2025video] and can capture subtle errors, temporal nuances, and intent—details that scalar or categorical labels often miss. Finally, a few recent skilled activity video-language datasets [burgess2025video, li2024egoexo, huang2024egoexolearn, grauman2024ego] use experts to obtain free-form language descriptions akin to verbal coach feedback but do not provide rigorous evaluation benchmarks or tasks, making it difficult to assess how well modern video models can understand nuanced physical human skills.
@@
To address these issues, we introduce ExAct, a video-language benchmark consisting of 3,521 video–question–answer (VQA) pairs designed to evaluate expert-level understanding of skilled physical human actions (sample questions shown in Figure 1). ExAct covers 11 activities across 6 diverse physical domains: Sports (Basketball, Soccer, Bouldering), Bike Repair, Cooking, Health (COVID‑19 Safety, CPR), Music (Guitar, Piano, Violin), and Dance, as shown in Figure 2. To construct ExAct, we build on the fine-grained expert commentaries derived from the Ego-Exo4D dataset [grauman2024ego]. The original Ego-Exo4D expert commentaries have several crucial limitations: 1) the original expert commentaries are unstructured and lengthy, 2) they contain automatic speech recognition (ASR) errors, and 3) they include redundant or irrelevant content that is not directly tied to the observed actions. Furthermore, evaluating the quality of open-ended captions/descriptions in the original form of Ego-Exo4D expert commentaries is inherently difficult using standard language metrics such as CIDEr [vedantam2015cider], BLEU [papineni2002bleu], or ROUGE [lin2004rouge], which do not accurately reflect the correctness or relevance of the instructional feedback in the context of skilled physical activities. To address these challenges, we first apply a structured annotation pipeline that rewrites the original commentaries into concise, self-contained feedback commentaries. Then, our proposed ExAct benchmark evaluates expert action understanding as a multiple-choice question-answering task, which uses a well-defined metric of question-answering accuracy. This multiple-choice question-answering format eliminates the ambiguity and reproducibility issues associated with open-ended evaluation by providing clearly defined answers and controllable question difficulty through the design of distractor options.
We evaluate several state-of-the-art VLMs, including Gemini 2.5 Pro [comanici2025gemini], GPT-4.1 [hurst2024gpt], GPT-4o [hurst2024gpt], Gemini 1.5 Pro [team2024gemini], LLaVA-Video [zhang2024video], LLaVA-OneVision [li2024llava], Qwen2.5-VL [bai2025qwen2], VideoLLaMA [zhang2025videollama], InternVL2.5 [chen2024internvl], and PerceptionLM [cho2025perceptionlm] on our new ExAct benchmark. Our results reveal that compared to human experts, most modern VLMs achieve poor results on ExAct. In particular, Gemini 2.5 Pro, the best-performing model in our experiments, achieves 55.35% accuracy. In comparison, non-expert humans achieve 61.86% accuracy, while domain experts achieve 82.02% accuracy. This substantial gap highlights the limitations of current VLMs in expert-level understanding of physical human skills. We hope that ExAct will serve as a rigorous evaluation benchmark for measuring expert-level understanding of skilled human actions, thus laying the foundation for AI systems that support enhanced human skill learning.
@@ 2 Related Work
Vision-Language Models (VLMs). Recent advances in Vision-Language Models (VLMs) have demonstrated impressive capabilities in understanding visual content. Models such as Gemini 2.5 Pro [comanici2025gemini], GPT-4.1 [hurst2024gpt], GPT-4o [hurst2024gpt], Gemini 1.5 [team2024gemini], LLaVA-OneVision [li2024llava], Qwen2.5-VL [bai2025qwen2], VideoLLaMA [zhang2025videollama], and InternVL2.5 [chen2024internvl] have achieved strong performance in tasks such as action recognition, video captioning, and visual question answering. Although these models show remarkable generalization, their outputs are often limited to high-level descriptions of the image/video content, rather than a detailed, fine-grained understanding of physical human actions and skills. More recent video-centric variants, such as LLaVA-Video [zhang2024video] and VideoLLaMA [zhang2025videollama], attempt to extend static image capabilities to temporal inputs. However, these models still struggle to understand the fine-grained nuances of complex physical human activities, which are essential for human skill understanding. Specifically, most existing VLMs cannot assess how well a task is performed, nor articulate the strengths of the execution and the areas requiring improvement [burgess2025video, cai2024temporalbench, pan2025basket]. To address this gap, our proposed ExAct provides a rigorous benchmark to evaluate expert-level understanding of physical human skills.
@@
Multimodal and VQA Benchmarks. Diverse multimodal benchmarks have emerged to evaluate VLM performance, particularly in the video domain. ActivityNet-QA [yu2019activitynet] assesses the temporal reasoning of activities within longer videos via question-answering. NExT-QA [xiao2021next] focuses on compositional temporal reasoning. STAR [wu2024star] emphasizes situated reasoning. Benchmarks such as PerceptionTest [patraucean2023perception] probe the perceptual abilities of VLMs in various modalities. TemporalBench [cai2024temporalbench] targets fine-grained temporal understanding. MV-Bench [li2024mvbench] assesses temporal comprehension across 20 challenging tasks. EgoSchema [mangalam2023egoschema] focuses on egocentric human actions. MMWorld [he2024mmworld] aims to evaluate embodied agents in simulated environments. MLVU [zhou2024mlvu], Video-MMMU [hu2025video], MMVU [zhao2025mmvu], Video-MMLU [song2025video], and Video-MME [fu2024video] aim to evaluate modern MLLMs for complex video question answering tasks. Finally, SEED-Bench [li2023seed] and MMBench [liu2024mmbench] evaluate various multimodal abilities of MLLMs. However, most existing benchmarks focus on factual recall (e.g., “What color is the car?”), event-level understanding (e.g., “What action is being performed?”), or basic temporal reasoning (e.g., “What happened before this?”), and are not designed to capture the subtle nuances of skilled human actions. In contrast, our newly proposed ExAct benchmark focuses explicitly on expert-level analysis of physical human skills.
@@
Video-based Skill Assessment Benchmarks. Several recent works have focused on developing methods to assess human skills from video. Benchmarks such as MITDive [pirsiavash2014assessing], UNLV-Dive [parmar2017learning], MTL-AQA [parmar2019and], and FineDiving [xu2022finediving] provide temporally segmented videos with action labels or action quality scoring for diving. FP-Basket [bertasius2017baller] and BASKET [pan2025basket]) focus on basketball, while LOGO [zhang2023logo]) provides human judgment scores for artistic swimming. Similar efforts also exist in other sports, including figure skating [pirsiavash2014assessing] and golf [mcnally2019golfdb]. Furthermore, datasets such as JIGSAWS [ahmidi2017dataset], BEST [Doughty_2019_CVPR], and EgoExoLearn [huang2024egoexolearn] focus on scenarios beyond sports, such as surgical tasks and daily activities. Most recently, the Ego-Exo4D [grauman2024ego] dataset introduces large-scale egocentric and exocentric video of skilled human activities. Ego-Exo4D includes spoken expert commentaries, offering a unique expert-level supervisory signal for understanding human skills. However, these expert commentaries are typically highly unstructured, noisy due to ASR errors, and often contain irrelevant information. Moreover, Ego-Exo4D does not provide a formal evaluation benchmark/task associated with such expert commentaries. In our work, we leverage such unstructured expert commentaries and construct a rigorous, expert-validated, and easy-to-evaluate ExAct benchmark enabling evaluation of expert-level understanding of physical human actions/skills.
@@ 3 ExAct Benchmark Construction
We construct ExAct using a four-stage pipeline. In stage I, we pre-process raw expert commentaries using GPT-4o, correcting ASR errors and segmenting them into concise, self-contained feedback commentaries. In Stage II, we construct multiple-choice QA pairs, each consisting of one correct commentary and four distractors. In Stage III, we remove low-quality or biased samples through length filtering and blind-LLMs. Finally, in Stage IV, domain experts review each QA pair to ensure visual grounding and linguistic accuracy. Our benchmark construction pipeline is illustrated in Figure 3.
@@ 3.2 Stage II: Question and Answer Generation
Albeit simple in format, our questions effectively probe multi-faceted reasoning abilities—spanning temporal, causal, spatial, and domain-specific understanding—beyond mere visual recognition (see supplementary material S4 for examples).
@@ 3.3 Stage III: Generated Question Answer Filtering
After constructing initial QA samples (Subsection 3.2), we apply several additional filtering steps to ensure high-quality and unbiased samples. Prior studies [cai2024matryoshka, yue2024mmmu] have shown that large language models (LLMs) can often exploit subtle statistical or stylistic patterns in multiple-choice answers to correctly identify the ground-truth option without relying on visual input. Such language-driven shortcuts, often referred to as language-related bias, pose a serious threat to the integrity of robust evaluation. To mitigate such biases, we adopt two filtering strategies:
1) Length similarity: Similar to human test-takers, LLMs may exploit surface-level cues such as answer length to eliminate implausible distractors. To mitigate this bias, we enforce a length similarity constraint: each distractor must be between 80% and 120% the length of the positive (i.e., correct) expert commentary, and the absolute word count difference must not exceed 8 words. QA samples that violate this constraint are excluded from the benchmark. Note that the 80–120% length range and 8-word absolute difference threshold were determined empirically through iterative pilot studies with human annotators.
2) Blind-LLM filtering: Inspired by findings from TemporalBench [cai2024temporalbench], we observe that LLMs can sometimes identify the correct answer by detecting shared linguistic patterns, particularly when distractors are lightly edited variants of the ground truth. To avoid such language-driven biases, we present each QA sample consisting solely of the five textual options without any video to 5 state-of-the-art LLMs (GPT-4o and DeepSeek-VL-R1 (proprietary), Qwen2.5-72B, LLaMA3.3-70B, and InternLM2.5-20B (open-source, all ranked highly on multiple LLM leaderboards)). Each model is prompted with the question: “Which expert commentary best matches the provided video?” If more than 20% of the models (i.e., exceeding the random chance of selecting the correct answer in a 5-way setup) select the positive expert commentary, we consider the sample susceptible to language-only bias and remove it from the dataset.